# Junctiond: Extending FaaS Runtimes with Kernel-Bypass

### 06 Mar 2024 v1

@@ Abstract
This report explores the use of kernel-bypass networking in FaaS runtimes and demonstrates how using Junction [9], a novel kernel-bypass system, as the backend for executing components in faasd can enhance performance and isolation. Junction achieves this by reducing network and compute overheads and minimizing interactions with the host operating system. Junctiond, the integration of Junction with faasd, reduces median and P99 latency by 37.33% and 63.42%, respectively, and can handle 5 times more throughput while decreasing latency by $2\times$2 × at the median and 3.5 times at the tail.
@@ 1 Introduction
Newer kernel-bypass systems, like Junction [9], bring the performance benefits without the complexities and resources overhead involved. In this work, we demonstrate how we can utilize these systems to increase throughput and reduce warm end-to-end latency in FaaS runtimes. Specifically, we evaluate how Junction [9] can be seamlessly integrated into faasd as its primary execution runtime, reducing tail latencies by up to 81% and increasing throughput by up to 5 times, without significantly increasing the number of allocated cores per server.
@@ 5 Evaluation
Tail latency vs load. Figure 6 shows the tail latency across varying request rates offered via the front-end load balancer. Junction can sustain $5\times$5 × more throughput while lowering the latency by $\sim 2\times$∼ 2 × at the median and $\sim 3.5\times$∼ 3.5 × at the tail. This reflects the compounding end-to-end benefit of using Junction across multiple components running in separate instances.
@@ 6 Conclusion
We discussed the use of kernel-bypass networking in Function as a Service (FaaS) runtimes, specifically using Junction as the backend for executing components in faasd. Our results show that using Junction can improve performance by reducing latency and increasing throughput, while also reducing the attack surface when compared to containers. Junction reduces both median and P99 latency by 37.33% and 63.42%, respectively, and can sustain $5\times$5 × more throughput while lowering the median latency by $2\times$2 × and the tail by $3.5\times$3.5 ×.

### 07 Mar 2024 v2

@@ Abstract
This report explores the use of kernel-bypass networking in FaaS runtimes and demonstrates how using Junction [9], a novel kernel-bypass system, as the backend for executing components in faasd can enhance performance and isolation. Junction achieves this by reducing network and compute overheads and minimizing interactions with the host operating system. Junctiond, the integration of Junction with faasd, reduces median and P99 latency by 37.33% and 63.42%, respectively, and can handle 10 times more throughput while decreasing latency by $2\times$2 × at the median and 3.5 times at the tail.
@@ 1 Introduction
Newer kernel-bypass systems, like Junction [9], bring the performance benefits without the complexities and resources overhead involved. In this work, we demonstrate how we can utilize these systems to increase throughput and reduce warm end-to-end latency in FaaS runtimes. Specifically, we evaluate how Junction [9] can be seamlessly integrated into faasd as its primary execution runtime, reducing tail latencies by up to 81% and increasing throughput by up to 10 times, without significantly increasing the number of allocated cores per server.
@@ 5 Evaluation
Tail latency vs load. Figure 6 shows the tail latency across varying request rates offered via the front-end load balancer. Junction can sustain $10\times$10 × more throughput while lowering the latency by $\sim 2\times$∼ 2 × at the median and $\sim 3.5\times$∼ 3.5 × at the tail. This reflects the compounding end-to-end benefit of using Junction across multiple components running in separate instances.
@@ 6 Conclusion
We discussed the use of kernel-bypass networking in Function as a Service (FaaS) runtimes, specifically using Junction as the backend for executing components in faasd. Our results show that using Junction can improve performance by reducing latency and increasing throughput, while also reducing the attack surface when compared to containers. Junction reduces both median and P99 latency by 37.33% and 63.42%, respectively, and can sustain $10\times$10 × more throughput while lowering the median latency by $2\times$2 × and the tail by $3.5\times$3.5 ×.